# The mechanism for creating“dynamical gravastar” black hole mimickers can also explain formation of “little red dots”

### v2

@@ I Introduction
In Section 4 we give a simplified model of a dynamical gravastar accreting matter from its environment, at a slow enough rate so that it grows by passing through the equilibrium states described in Section 2. Following this growth using the Mathematica notebook for dynamical gravastars shows that as the mimicker mass grows, the depth of the gravitational potential well in which latent energy is released correspondingly grows, and so does the red shift of exiting radiation. This calculation also shows that as the mimicker grows, it continues to obey the mathematical black hole radius–mass relation $R=2GM$ (which in geometrized units with $G=1$ as used in the Mathematica notebook is simply $R=2M$).
@@ IV An alternative interpretation based on the structure of dynamical gravastar black hole mimickers
In this section we propose an alternative mechanism for the formation of red dots. Like the “black hole star” proposal, it is a variant on conventional black hole ideas. But the redness in our proposal results not from absorption of energetic radiation by an opaque surround with subsequent re-emission at red wavelengths, but rather from the gravitational redshift of initially energetic radiation as it climbs out of the deep gravitational potential well associated with the gravastar. Moreover, in our proposal the source of the initial energetic radiation is not the violent accretion of matter by a central black hole, but rather the latent energy released in the first order phase transition to negative matter energy density at pressures above pjump, as postulated in the dynamical gravastar model. A brief sketch of this mechanism for “dot” formation was given in a post-publication Appendix C added to arXiv:2504.18690v4. We elaborate on our sketch in this section by studying a simple model for dynamical gravastar growth.
@@
Sample results of this study are given in Table I, and in Figs. 2 and 3. The columns of Table I, reading across, give the central pressure $p(0)$, the mass $M$ as seen at effective infinity, the radius ${\rm rjump}=r({\rm pjump})$ where the energy density jumps, the radius $r_{\rm MIN}$ at the kink minima seen in Figs. 2 and 3 which plot 12m(r)/r1-2m(r)/r, the central value of $\log g_{00}(0)=\nu(0)$, and finally the jump value of $\log g_{00}({\rm rjump})=\nu({\rm rjump})$. Scanning the table, one sees that (i) the mass $M$ increases monotonically, and very nonlinearly, with $p(0)$, (ii) the gravastar nominal horizon at the kink minimum obeys $r_{\rm MIN}\simeq 2M$, and (iii) the central and jump radius values of $\log g_{00}(r)=\nu(r)$ decrease monotonically with increasing $M$, becoming very large negative for large $M$. From Figs. 2 and 3 we see that ${\rm denom}=1-2m(r)/r$ always has a clear minimum defining $r_{\rm MIN}$. Also, when $p(0)$ is not too close to ${\rm pjump}=1$, there is a sharp transition region from the exterior to the interior structure, with most of the transition taking place at radii above the jump radius, as already seen in the stacked plots of Fig. 1. This means that as the mass $M$ increases, the latent energy released inside the jump radius has an increasingly large redshift to exit to infinity, giving the observational presentation of an energetic “red dot”.
@@ V Observational questions
We note two questions that may be addressable in future observations. First, is there an observed correlation between the dot radii and masses, and if so, does this approximate the geometrized black hole radius–mass relation $R=2M$? Second, is there a way of distinguishing observationally between redness arising from absorption of radiation by an opaque layer with subsequent thermalized re-radiation, and redness arising as a gravitational redshift effect?
@@ Appendix A Mathematica code for inversion of $p(r)$
Figs. 2 and 3 can be plotted with the Mathematica code given in Appendix A of adler1 . Table I uses the following additional lines of code to invert the interpolating function $p(r)$ and then calculate rjump:

### v3

@@ I Introduction
In Section 4 we give a simplified model of a dynamical gravastar accreting matter from its environment, at a slow enough rate so that it grows by passing through the equilibrium states described in Section 2. Following this growth using the Mathematica notebook for dynamical gravastars shows that as the mimicker mass grows, the depth of the gravitational potential well in which latent energy is released correspondingly grows, and so does the red shift of exiting radiation. As the mimicker grows, it continues to obey the mathematical black hole radius–mass relation (in geometrized units) $R\simeq 2M$, and has a photosphere with asymptotic pressure $p(r)\propto{\rm constant}\times e^{4M/r}$.
@@ IV An alternative interpretation based on the structure of dynamical gravastar black hole mimickers
In this section we propose an alternative mechanism for the formation of red dots. Like the “black hole star” proposal, it is a variant on conventional black hole ideas. But the redness in our proposal results not from absorption of energetic radiation by an opaque surround with subsequent re-emission at red wavelengths, but rather from the gravitational redshift of initially energetic radiation as it climbs out of the deep gravitational potential well associated with the gravastar. Moreover, in our proposal the source of the initial energetic radiation is not the accretion of matter by a central black hole, but rather the latent energy released in the first order phase transition to negative matter energy density at pressures above pjump, as postulated in the dynamical gravastar model. A brief sketch of this mechanism for “dot” formation was given in a post-publication Appendix C added to arXiv:2504.18690v4. We elaborate on our sketch in this section by studying a simple model for dynamical gravastar growth.
@@
Sample results of this study are given in Table I, and in Fig. 2. The columns of Table I, reading across, give the central pressure $p(0)$, the mass $M$ as seen at effective infinity, the radius ${\rm rjump}=r({\rm pjump})$ where the energy density jumps, the radius $r_{\rm MIN}$ at the kink minima seen in Fig. 2 which plot 12m(r)/r1-2m(r)/r, the central value of $\log g_{00}(0)=\nu(0)$, and finally the jump value of $\log g_{00}({\rm rjump})=\nu({\rm rjump})$. Scanning the table, one sees that (i) the mass $M$ increases monotonically, and very nonlinearly, with $p(0)$, (ii) the gravastar nominal horizon at the kink minimum obeys $r_{\rm MIN}\simeq 2M$, and (iii) the central and jump radius values of $\log g_{00}(r)=\nu(r)$ decrease monotonically with increasing $M$, becoming very large negative for large $M$. From Fig. 2 we see that ${\rm denom}=1-2m(r)/r$ always has a clear minimum defining $r_{\rm MIN}$. Also, when $p(0)$ is not too close to ${\rm pjump}=1$, there is a sharp transition region from the exterior to the interior structure, with most of the transition taking place at radii above the jump radius, as already seen in the stacked plots of Fig. 1. This means that as the mass $M$ increases, the latent energy released inside the jump radius has an increasingly large redshift to exit to infinity, giving the observational presentation of an energetic “red dot”.
The pressure distribution in a dynamical gravastar has a tail extending to infinity, and the radius where the optical thickness of this tail is unity constitutes the “photosphere”, giving the observed radius. An exact integration of the exterior region equations gives the formula
$p(r)={\rm pjump}e^{2(\nu(\rm{rjump})-\nu(r))}=p(\infty)e^{-2\nu(r)}~~~,$ (6)
and substituting the approximation (good except very near the jump radius) $g_{00}(r)\simeq 1-2M/r$ gives as a a good large radius fit to the tail (as illustrated in Fig. 3)
$p(r)={\rm pjump}e^{2(\nu(\rm{rjump})+2M/r)}=p(\infty)e^{4M/r}~~~.$ (7)
A priori, one does not know the conversion from units of the Mathematica notebook to physical units, which would be needed to give a prediction of the observed dot size.
@@ V Observational questions
We note two questions that may be addressable in future observations. First, is there an observed correlation between the dot radii and masses, and if so, can the luminosity tail be fitted with a simple functional form as in Eq. (7)? Second, is there a way of distinguishing observationally between redness arising from absorption of radiation by an opaque layer with subsequent thermalized re-radiation, and redness arising as a gravitational redshift effect?
@@
VII Acknowledgements
I wish to than Scott Tremaine for email correspondence emphasizing the role of the photosphere in determining the observed dot radius.
@@ Appendix A Mathematica code for inversion of $p(r)$
Fig. 2 can be plotted with the Mathematica code given in Appendix A of adler1 . Table I uses the following additional lines of code to invert the interpolating function $p(r)$ and then calculate rjump: